A rectangular garden has a perimeter of 188 feet. The width is 5 ft less than twice the length. Find the length and the width of the garden.
Answer by mananth(16949)

A rectangular garden has a perimeter of 188 feet. The width is 5 ft less than twice the length. Find the length and the width of the garden.
The perimeter is 188 feet
let the length be x
width will be 2x-5
perimeter = 2* length + 2 * width
188= 2x + 2*(2x-5)
188= 2x+4x -10
188+10= 6x
198/6= x
x=33 feet which is the length
Width will be 2x-5 = 2*33 - 5 = 66-5 = 61
Width = 61